OpenStudy (anonymous):

SOLVE: l x + 2 l = 5

OpenStudy (anonymous):

solve two equations, get two different answers one equation is \[x+2=5\] and the other is \[x+2=-5\]

OpenStudy (anonymous):

@satellite73 can you further explain please?

OpenStudy (anonymous):

explain how to solve \(x+2=5\) ?

OpenStudy (anonymous):

or explain how i got those two equations?

OpenStudy (anonymous):

@satellite73 how you got those two equations

OpenStudy (anonymous):

there are exactly two numbers whose absolute value is \(5\) one is \(5\) itself, in which case \(x+2=5\) the other is \(-5\) because \(|-5|=5\) and therefor it could be \[x+2=-5\]

OpenStudy (anonymous):

@satellite73 thank you would you mind answering 3 more?

OpenStudy (anonymous):

@satellite73 l x l > 5